# Modelling the evolution and influence of dust in cosmological simulations that include the cold phase of the interstellar medium

### v1

@@ 2 Modelling the formation & evolution of Dust
Here we describe the constituent elements of our dust model, including the assumed properties of grains and the evolutionary processes of grain seeding/nucleation, grain growth, grain destruction and grain size transfer. Particular focus is placed on aspects of the modelling that differ from previous studies. Our dust model is designed for compatibility with SWIFT SPH simulations (Schaller et al., 2016, 2024), particularly the upcoming COLIBRE simulations. All the processes detailed below (seeding, growth and destruction of grains), are implemented on a particle-by-particle basis, using the properties of a given gas resolution element. Throughout, we characterise the dust content in terms of the mass fractions, either as the dust-to-gas ratio ($\mathcal{DTG}$) or dust-to-metal ratio ($\mathcal{DTZ}$). With the exception of direct destruction by supernovae (SNe) or astration, all our processes depend on gas density. For dense-gas processes (such as grain growth), we consider an effective density indicated using (i.e. $\rho^{\prime}$, $n^{\prime}_{\rm H}$), which is related to the physical density by a clumping factor, C, as detailed in §3.7. The total element abundances include both the abundances of metals that remain in the gas-phase, which we refer to as the diffuse component, and the abundance that is depleted into dust grains. Throughout, we use the fiducial COLIBRE stellar initial mass function (IMF) of Chabrier 2003, with a range $\in[0.1,100]\;{\rm M}_{\odot}$ (hereafter C03).
@@ 2.1 Types of dust grains
Conversely, silicates are generally heteronuclear, with a variety of different elements contributing significantly to their mass. Strong candidate minerals from stoichiometry are olivine, pyroxene and intermediate/composite grains (e.g. Fogerty et al., 2016). We explore both olivine and pyroxene mineral series, adopting olivine as our fiducial silicate. We then treat Mg and Fe end-members as distinct subspecies tracked independently: forsterite (Mg2SiO4) and fayalite (Fe2SiO4) in the case of olivine [^1] Enstatite (Mg2Si2O6) and ferrosilite (Fe2Si2O6) for pyroxenes.. While we limit the silicates to a particular mineral series, the choice to track Mg and Fe end-member silicates separately allows the ratio of these species to freely float given the local chemical abundances, such that they are not artificially limited by a fixed ratio of the two. In particular, the ratio of Mg to Fe species can then reflect variations in $\alpha$-enrichment, $[\rm{\alpha/Fe}]$.
@@
Silicate (Olivine) Fayalite, Forsterite FeMgSiO4 171.85 3.5×$\times 10^{-4}$ Si | Fe | Mg 28.09 | 55.85 | 24.31 99.3
A common simplifying assumption for grain structure is that of spherical grains. This allows a grain’s structure to be characterised by a single value, its radius $r_{\rm grain}$. While this greatly simplifies the equations for dust evolution, the assumption of sphericity has certain limitations. Some degree of asphericity in cosmic dust is evident in harvested samples (e.g. the STARDUST mission, Brownlee, 2014), and by the polarisation of light in the ISM (e.g. Cho & Lazarian, 2007). While the degree of asphericity across interstellar environments is hard to constrain, the assumption of spherical grains is extreme in that it minimises their surface-to-volume ratio, with implications for their interaction with gas or other grains. However, sphericity is degenerate with other physical properties modulating the growth/destruction rates of dust, which are themselves typically empirically determined or calibrated. For our purposes, the assumption of an effective spherical grain is adopted for simplicity.
@@ 2.2.2 SN grain production
In addition to synthesising large quantities of metals, the shocks from SNe are thought to facilitate significant dust grain formation. Local SN remnants provide valuable case studies for these processes (e.g. Dunne et al., 2009). However, the overall dust yields from SNe are particularly uncertain, as strong reverse shocks can also destroy grains (section 2.4). At very early times, core-collapse SNe are the primary candidates for dust grain formation (e.g. Tielens, 1998; Dwek, 1998). Observations of $z\approx 7$ radio galaxies and quasars with $M_{\rm dust}>10^{7}\;{\rm M_{\odot}}$ have been taken to suggest that SN on average produce $\approx 1\;{\rm M_{\odot}}$ of dust (Valiante et al., 2009, 2011; Watson et al., 2015), while other observations of low-luminosity star-forming systems at $z>6.5$ (Bouwens et al., 2012) infer very low levels of dust attenuation.
It is less clear whether type Ia SNe (SNIa) net produce dust (e.g. Kozasa et al., 2009; Nozawa et al., 2011). Some studies have proposed SNIa as a viable dust production mechanism (Clayton et al., 1997; Travaglio et al., 1999). However, observations of local SNIa remnants do not show significant evidence of new dust formation (Gomez et al., 2012). Nozawa et al. (2011) posits that this lack of evidence may be explained by a lower formation radius and subsequent rapid destruction in SNIa, preventing them from being an appreciable source of ISM dust grains. Given these findings, we omit dust production by SNIa, and consider only core-collapse supernova (CCSN) production.
@@ 2.2.4 Size distribution of seeded grains
In addition to seeding the mass in each chemical species of grain, the initial size distribution of the seeded grains must also be set. We assume that seed grains are dominated by larger size ($a\approx 0.1$ µm). This reflects studies of AGB winds (Groenewegen, 1997; Yasuda & Kozasa, 2012; Asano et al., 2013b), and is consistent with the inference of $r_{\rm grain}\gtrsim 0.01$ µm in core-collapse SNa (e.g. Nozawa et al., 2007). We initiate the size distribution with 90% of the mass in large grains and the remaining 10% in small grains. This is an order-of magnitude estimate, chosen to be between the values seen in AGB winds (e.g. Asano et al., 2013a) and the observed Milky-Way (MW) mass distribution (e.g. Mathis et al., 1977). We note that the size distribution of seed grains, like their absolute masses, has little effect on the resultant dust content in galaxies in our model, as this is set by the growth of grains and its equilibration with destructive processes in the ISM (see §4.1).
@@ 2.2.5 Comparing dust yields from seeding channels
In Fig. 1 we show the dust yields (top row) and returned dust-to-metal ratios (bottom row) for CCSN (left column) and AGB (right column) as functions of their ZAMS mass, $M_{\rm init}$. A striking feature is that the yields advocated by Dwek (1998) are higher than those we use by about an order of magnitude. In addition, the dust-to-metal ratio of returned material in the Dwek (1998) model is of the same order as that seen in the local ISM ($\approx 0.3$, Inoue, 2003). This difference is profound as it implies different roles for evolutionary processes in reaching the dust content of galaxies like the Milky Way. Our model requires dust masses to grow by a factor $\sim 10$ in the ISM, while the Dwek (1998) yields require either no mass evolution, or an equilibrium of destruction and growth mechanisms. In particular, our lower level of seeding requires modelling of the cold, dense ISM conducive to efficient grain growth (see §2.3), as opposed to a level of seeding requiring only maintenance. This difference is discussed further in the following section.
@@ 2.3 Grain growth
where ${n^{\prime}_{\rm H}}$ is the effective local number density of hydrogen, $T$ is the local gas temperature, $S_{\rm acc}$ is the sticking probability for accretion, $\tau_{\rm G}$ is the accretion timescale normalisation for a given grain species, and $\epsilon_{j}$ and $\epsilon_{j,\odot}$ represent the local and solar abundance of the bottleneck abundance in the diffuse (gas) phase relative to hydrogen, respectively. For our purposes, we assume a fixed $S_{\rm acc}$ value of 0.3, with $\tau_{\rm G}$ for silicate and carbon grains taken from Hirashita & Voshchinnikov (2013) and listed in Table 1. The fractional change in the mass of a given dust species due to accretion is computed over a timestep of $\Delta t$ for each gas particle, as $\exp{(\Delta t/\tau_{\rm acc})}$.
@@
This is reflective of the fact that total accretion in the silicate species is freely divisible between Mg and Fe endmembers, such that the total abundance of Mg and Fe can be considered a single reservoir of diffuse metals available for accretion. With O and Si always present in silicates with a fixed abundance ratio, these elements can be treated as independent reservoirs.
@@
This accretion for the overall silicate species is computed and the total accreted mass is divided between endmember species proportionally to the relative abundance ratio of diffuse endmember elements in the ISM.
@@
The primary sites of accretion are predicted to be in molecular cloud environments, with local densities of $n_{\rm H}\gtrsim 10^{3}\;{\rm cm}^{-3}$ (Hirashita, 2000), where $n_{\rm H}$ is the hydrogen number density where $n_{\rm H}=m_{\rm p}^{-1}{\rm X_{g}}\rho$ for proton mass $m_{\rm p}$ and hydrogen mass fraction, ${\rm X_{g}}$. This is supported by the enhanced depletion observed in dense clouds (Savage & Sembach, 1996). For this reason, to properly account for accretion it is important that molecular clouds are represented. Zhukovska et al. (2008) attribute smaller dust nucleation rates compared to Dwek (1998) from stellar sources to a more explicit treatment of dust growth in molecular clouds. A conflation of stellar nucleation and accretion could be one explanation for the significantly higher dust yields displayed for Dwek (1998) in Fig. 1.
@@
Using dust yields that implicitly include molecular cloud grain growth may be a useful strategy for simulations that do not attempt to model such dense gas structures (McKinnon et al., 2017; Davé et al., 2019; Vijayan et al., 2019). However, our model has been developed for galaxy formation simulations that include multiphase gas and that can attempt to distinguish molecular clouds from the diffuse ISM, such that using boosted yields would effectively ‘double count’ accretion. Despite the relatively high densities reached in our simulations, we are still far from resolving the dense cores of molecular clouds where these grains are thought to accrete efficiently. As a result we apply a clumping factor, $C$, to boost the densities input into the accretion rate. Our fiducial scheme uses a variable boost factor that rises monotonically from $C=1$ to 100 between $-1<\log_{10}{n_{\rm H}/{\rm cm^{-3}}}<2$ and is clipped to the boundary values on either side. This is found to reproduce observed dust-to-metal ratios within gas at ISM densities ($n_{\rm H}\gtrsim 0.1\;{\rm cm^{-3}}$). We investigate the clumping factor further in §3.7.
@@ 2.4 Grain destruction
2.4.1 Sputtering
Sputtering is the erosion of dust grains through high-velocity interactions with gas particles in hot gas. We use the sputtering prescription of Tsai & Mathews (1995), which uses the grain shrinkage rate
@@ 2.4.1 Sputtering
where $T_{0}$ is the temperature normalisation, set to $2\times 10^{6}$ K. We note that the sputtering timescale is highly sensitive to the temperature, such that dust in gas heated to temperatures $T>T_{0}$ will be destroyed very quickly. Note that as a diffuse-gas process, we use the unmodulated density, $n_{\rm H}$, for shattering (as opposed to $n_{\rm H}^{\prime}$, modulated by a clumping factor, $C$).
@@ 2.4.2 Astration
Another process to consider is the destruction of grains upon their reincorporation into stellar atmospheres, also known as astration. In our model, the dust mass that is present in a gas particle as it is being converted to a stellar population is transferred back to the diffuse element mass fraction array stored in each particle, in the proportions dictated by each grain compound. These diffuse elemental mass fractions then come to represent the total elemental mass fractions for the resultant star particle.
@@ 2.4.3 SN shocks
While the effect of SN shocks on grains may be considered an extension of sputtering, the shocks of individual SNe are not resolved in galaxy-scale simulations. Instead, in COLIBRE bulk feedback and star formation events are modelled stochastically (§3.3, §3.2). For consistency, we implement a stochastic approach linked directly to the feedback module - dust is totally destroyed in the particles stochastically selected to receive SN feedback [^3] For either thermal or kinetic implementations of SN feedback.. This is handled in the same way as astration - on being hit, mass in dust is redistributed into the diffuse phase of the constituent elements, and dust abundances are set to zero, all within the struck particle.

### v2

@@ 3 Modelling the formation & evolution of Dust
We now describe the constituent elements of our dust model, including the assumed properties of grains and the evolutionary processes of grain seeding/nucleation, grain growth, grain destruction and grain size transfer. Particular focus is placed on aspects of the modelling that differ from previous studies. Our dust model is designed for compatibility with SWIFT SPH simulations (Schaller et al., 2016, 2024), particularly the upcoming COLIBRE simulations, as detailed above. All the processes detailed below (seeding, growth and destruction of grains), are implemented on a particle-by-particle basis, using the properties of a given gas resolution element. Throughout, we characterise the dust content in terms of the mass fractions, either as the dust-to-gas ratio ($\mathcal{DTG}$) or dust-to-metal ratio ($\mathcal{DTZ}$). With the exception of direct destruction by supernovae (SNe) or astration, all our processes depend on gas density. For dense-gas processes (such as grain growth), we consider an effective density indicated using (i.e. $\rho^{\prime}$, $n^{\prime}_{\rm H}$), which is related to the physical density by a clumping factor, C, as detailed in §3.8. The total element abundances include both the abundances of metals that remain in the gas-phase, which we refer to as the gas-phase component, and the abundance that is depleted into dust grains. Throughout, we use the fiducial COLIBRE stellar initial mass function (IMF) of Chabrier 2003, with a range $\in[0.1,100]\;{\rm M}_{\odot}$ (hereafter C03).
@@ 3.1 Types of dust grains
Conversely, silicates are generally heteronuclear, with a variety of different elements contributing significantly to their mass. Strong candidate minerals from stoichiometry are olivine, pyroxene and intermediate/composite grains (e.g. Fogerty et al., 2016). We explore both olivine and pyroxene mineral series, adopting olivine as our fiducial silicate. We then treat Mg and Fe end-members as distinct subspecies tracked independently: forsterite (Mg2SiO4) and fayalite (Fe2SiO4) in the case of olivine [^9] Enstatite (MgSiO3) and ferrosilite (FeSiO3) for pyroxenes.. While we limit the silicates to a particular mineral series, the choice to track Mg and Fe end-member silicates separately allows the ratio of these species to freely float given the local chemical abundances, such that they are not artificially limited by a fixed ratio of the two. In particular, the ratio of Mg to Fe species can then reflect variations in $\alpha$-enrichment, $\rm{\alpha/Fe}]$.
@@
[…]
A common simplifying assumption for grain structure is that of spherical grains. This allows a grain’s structure to be characterised by a single value, its radius $r_{\rm grain}$. While this greatly simplifies the equations for dust evolution, the assumption of sphericity has certain limitations. Some degree of asphericity in cosmic dust is evident in harvested samples (e.g. the STARDUST mission, Brownlee, 2014), and by the polarisation of light in the ISM (e.g. Cho & Lazarian, 2007). In addition, grains may have a degree of porosity, increasing their surface area and collisional and optical properties (Hirashita & Il’in, 2022). While the degree of asphericity and porosity across interstellar environments is hard to constrain, the assumption of spherical, aporous grains is extreme in that it minimises their surface-to-volume ratio, with implications for their interaction with gas or other grains. However, sphericity is degenerate with other physical properties modulating the growth/destruction rates of dust, which are themselves typically empirically determined or calibrated. For our purposes, the assumption of an effective spherical grain is adopted for simplicity.
@@ 3.2.2 SN grain production
In addition to synthesising large quantities of metals, the shocks from SNe are thought to facilitate significant dust grain formation. Local SN remnants provide valuable case studies for these processes (e.g. Dunne et al., 2009; Shahbandeh et al., 2023; Milisavljevic et al., 2024). However, the overall dust yields from SNe are particularly uncertain, as strong reverse shocks can also destroy grains (e.g. Kirchschlager et al. 2019; Kirchschlager et al. 2023, see § 3.4). At very early times, core-collapse SNe are the primary candidates for dust grain formation (e.g. Tielens, 1998; Dwek, 1998). Observations of $z\approx 7$ radio galaxies and quasars with $M_{\rm dust}>10^{7}\;{\rm M_{\odot}}$ have been taken to suggest that SN on average produce $\approx 1\;{\rm M_{\odot}}$ of dust (Valiante et al., 2009, 2011; Watson et al., 2015), while populations of galaxies exhibiting much lower dust-to-metal ratios have also been observed down to $z\approx 4$ (Burgarella et al., 2025).
It is less clear whether type Ia SNe (SNIa) net produce dust (e.g. Kozasa et al., 2009; Nozawa et al., 2011). Some studies have proposed SNIa as a viable dust production mechanism (Clayton et al., 1997; Travaglio et al., 1999). However, observations of local SNIa remnants do not generally show evidence of significant net dust formation (e.g. Gomez et al., 2012), though evidence of dust formation in a SNIa remnant has been observed by Wang et al. (2024). Nozawa et al. (2011) posits that this lack of evidence may be explained by a lower formation radius and subsequent rapid destruction in SNIa, preventing them from being an appreciable source of ISM dust grains. Given these findings, we omit dust production by SNIa, and consider only core-collapse supernova (CCSN) production.
@@ 3.2.4 Size distribution of seeded grains
In addition to seeding the mass in each chemical species of grain, the initial size distribution of the seeded grains must also be set. We assume that seed grains are dominated by larger size ($a\approx 0.1$ µm). This reflects studies of AGB winds (Groenewegen, 1997; Yasuda & Kozasa, 2012; Asano et al., 2013b), and is consistent with the inference of $r_{\rm grain}\gtrsim 0.01$ µm in core-collapse SNa (e.g. Nozawa et al., 2007). We initiate the size distribution with 90% of the mass in large grains and the remaining 10% in small grains. This is an order-of magnitude estimate, chosen to be between the values seen in AGB winds (e.g. Asano et al., 2013a) and the observed Milky-Way (MW) mass distribution (e.g. Mathis et al., 1977). We note that the size distribution of seed grains, like their absolute masses, have little effect on the dust content of low-redshift galaxies that we present for out model, as this is set by the growth of grains and its equilibration with destructive processes in the ISM (see §4.2). However at higher redshift, or especially rapidly forming galaxies, the yield contribution can be more significant, with dust produced through yields before evolutionary processes can equilibrate [^10] In section 4.3, we show that in our model the total contribution from direct stellar yields is subdominiant to growth processes for $z\lesssim 11$ (§ 4.3, Fig. 7)..
@@ 3.2.5 Comparing dust yields from seeding channels
In Fig. 1 we show the dust yields (top row) and returned dust-to-metal ratios (bottom row) for CCSN (left column) and AGB (right column) as functions of their ZAMS mass, $M_{\rm init}$. A striking feature is that the yields advocated by Dwek (1998) are higher than those we use by about an order of magnitude. In addition, the dust-to-metal ratio of returned material in the Dwek (1998) model is of the same order as that seen in the local ISM ($\approx 0.3$, Inoue, 2003). This difference is profound as it implies different roles for evolutionary processes in reaching the dust content of galaxies like the Milky Way. Our model requires dust masses to grow by a factor $\sim 10$ in the ISM, while the Dwek (1998) yields require either no mass evolution, or an equilibrium of destruction and growth mechanisms. In particular, our lower level of seeding requires modelling of the cold, dense ISM conducive to efficient grain growth (see §3.3), as opposed to a level of seeding requiring only maintenance. We note that some studies use yields of order 10% but still require growth to reproduce observables (e.g. Lower et al., 2024; Choban et al., 2025).
While the Dwek (1998) yields are empirically motivated, with some arbitrary choices with regards to grain nucleation properties, some more recent observations of post-shock region of supernovae support this high-yield paradigm (e.g. Shahbandeh et al., 2023). The Zhukovska et al. (2008) yields are intended to be decoupled from subsequent moledcular cloud growth, and are consistent with the limited measurements of presolar grains in the solar system, as well as the lower end of supernovae observations. Ultimately, these differences showcase the uncertainty in the direct yields of grains and differential role of growth in dust models, explored further in the following section.
@@ 3.3 Grain growth
where $n^{\prime}_{\rm H}$} is the effective local number density of hydrogen, $T$ is the local gas temperature, $S_{\rm acc}$ is the sticking probability for accretion, $\tau_{\rm G}$ is the accretion timescale normalisation for a given grain species, and $\epsilon_{j}$ and $\epsilon_{j,\odot}$ represent the local and solar abundance of the bottleneck abundance in the gas-phase relative to hydrogen, respectively. For our purposes, we assume a fixed $S_{\rm acc}$ value of 0.3, with $\tau_{\rm G}$ for silicate and carbon grains taken from Hirashita & Voshchinnikov (2013) and listed in Table 2. The fractional change in the mass of a given dust species due to accretion is computed over a timestep of $\Delta t$ for each gas particle, as $\exp{(\Delta t/\tau_{\rm acc})}$.
@@
This is reflective of the fact that total accretion in the silicate species is freely divisible between Mg and Fe endmembers, such that the total abundance of Mg and Fe can be considered a single reservoir of gas-phase metals available for accretion. With O and Si always present in silicates with a fixed abundance ratio, these elements can be treated as independent reservoirs.
@@
This accretion for the overall silicate species is computed and the total accreted mass is divided between endmember species proportionally to the relative abundance ratio of gas-phase endmember elements in the ISM.
@@
Accretion proceeds most efficiently in cold, dense environments where dust constituent elements are available in the gas-phase. Accounting for how dust is distributed, the majority of accretion is posited to take place in the cold, neutral medium (e.g. Jenkins, 2009; Hirashita, 2012). This represents local densities of $n_{\rm H}\gtrsim 10\;{\rm cm}^{-3}$ (Hirashita, 2000), where $n_{\rm H}$ is the hydrogen number density where $n_{\rm H}=m_{\rm p}^{-1}{\rm X_{g}}\rho$ for proton mass $m_{\rm p}$ and hydrogen mass fraction, $\rm X_{g}$}. Efficient accretion in these contexts is supported by the enhanced depletion observed in dense clouds (Savage & Sembach, 1996). For this reason, to properly account for accretion it is important that dense, molecular media are represented. Zhukovska et al. (2008) attribute smaller dust nucleation rates compared to Dwek (1998) from stellar sources to a more explicit treatment of dust growth in molecular clouds. A conflation of stellar nucleation and accretion could be one explanation for the significantly higher dust yields displayed for Dwek (1998) in Fig. 1, though higher dust yields now observed in individual CCSNe may support the higher level grain seeding paradigm (Shahbandeh et al., 2023) comparable to Dwek (1998).
@@
Dust yields that implicitly include some molecular cloud grain growth can be useful for simulations that do not attempt to model such dense gas structures (McKinnon et al., 2017; Davé et al., 2019; Vijayan et al., 2019). However, our model has been developed for galaxy formation simulations that include multiphase gas and that can attempt (at sufficient resolution) to distinguish molecular gas clouds from the diffuse ISM, such that using yields that implicitly include an accreted component could effectively ‘double count’ accretion. Despite the relatively high densities reached in our simulations, we are still far from resolving the dense cores of molecular clouds where these grains are thought to accrete efficiently. As a result we apply a clumping factor, $C$, to boost the densities input into the accretion rate. Our fiducial scheme uses a variable boost factor that rises monotonically from $C=1$ to 100 between $1<\log_{10}{n_{\rm H}/{\rm cm^{-3}}}<2$ and is clipped to the boundary values on either side. This is found to reproduce observed dust-to-metal ratios within gas at ISM densities ($n_{\rm H}\gtrsim 0.1\;{\rm cm^{-3}}$). This provides a means to calibrate across regimes in resolution. We investigate the clumping factor further in §3.8. For our fiducial clumping with representative properties of the cold, neutral medium ($T=100$ K, $n_{\rm H}=10\;{\rm cm}^{-3}$) at $0.1\;Z_{\odot}$ a small carbon (silicate) grain in an environment where 5% of the bottleneck element is available for depletion in the gas-phase gives $\tau_{\rm acc}=26.4$ Myr ($\tau_{\rm acc}=14.6$ Myr), of comparable order to the ‘enhanced’ accretion model found necessary by Choban et al. (2025).
@@ 3.4 Grain destruction
3.4.1 Thermal Sputtering
Thermal sputtering is the erosion of dust grains through high-velocity interactions with gas particles in hot gas. We use the thermal sputtering prescription of Tsai & Mathews (1995), which uses the grain shrinkage rate
@@ 3.4.1 Thermal Sputtering
where $T_{0}$ is the temperature normalisation, set to $2\times 10^{6}$ K. We note that the sputtering timescale is highly sensitive to the temperature, such that dust in gas heated to temperatures $T>T_{0}$ will be destroyed very quickly at ISM densities. In hot gas at the lowst CGM densities the timescale can become long again. Note that because thermal sputtering occurs in the diffuse-gas phase, we use the unmodulated density, $n_{\rm H}$, for sputtering (as opposed to $n_{\rm H}^{\prime}$, modulated by a clumping factor, $C$).
@@ 3.4.2 Astration
Another process to consider is the destruction of grains upon their reincorporation into stellar atmospheres, also known as astration. In our model, the dust mass that is present in a gas particle as it is being converted to a stellar population is transferred back to the gas-phase element mass fraction array stored in each particle, in the proportions dictated by each grain compound. These gas-phase elemental mass fractions then come to represent the total elemental mass fractions for the resultant star particle.
@@ 3.4.3 SN shocks
The effect of SN shocks on grains may be considered a complement to the thermal sputtering we implement to account for resolved heating processes. The shocks of individual SNe are not resolved in galaxy-scale simulations. Instead, in COLIBRE bulk feedback and star formation events are modelled stochastically (§2.2, §2.1). For consistency, we implement a stochastic approach linked directly to the feedback module - dust is totally destroyed in the particles stochastically selected to receive SN feedback [^12] For either thermal or kinetic implementations of SN feedback.. This is handled in the same way as astration - on being hit, mass in dust is redistributed into the gas-phase of the constituent elements, and dust abundances are set to zero, all within the struck particle. The practical, stochastic implementation of this approach paired with the complex varying properties of the gas receiving feedback make it difficult to compute a timescale from first principles. To estimate the timescale for dust destruction by SNe in our simulations, we compare the average feedback-hit particles in the last 100 Myr (before $z=0$) with the $z=0$ dust properties in our fiducial model (see § 2.5), the ratio of which provides a cosmic timescale of around 420 Myr.